Use of ”mhequ.sty” (v1.5)

A simple labelled equation:

∑_{i=1}^{5} X_{i}^{j}X_{i}^{j} = y^{j} .  (1)

The name of the label at the right of the equation is generated by the draft option of the package. Without label, one gets the same:

∑_{i=1}^{5} X_{i}^{j}X_{i}^{j} = y^{j} .

A simple list of equations can be displayed either with one number per equation

f(x) = sin(x) + 1 ,  (2)
h(x) = f(x) + g(x) − 3 ,  (3)

or with one number for the whole list

\begin{aligned}f(x) & = sin(x) + 1 , \\ h(x) & = f(x) + g(x) − 3 . \end{aligned}  (4)

Of course, it can have no numbering at all.

f(x) = sin(x) + 1 ,
h(x) = f(x) + g(x) − 3 ,

Let us make a first group:

f(x) = sin(x) + 1 ,  (5a)
g(x) = cos(x) − x^{2} + 4 ,  (5b)
h(x) = f(x) + g(x) − 3 .  (5c)

You can refer to the whole block (5) or to one line, like (5a) for example. You can use any tag you like with the \tag command

x = y .  (tag)

Of course, you can also refer to (tag). You can use the usual equation environment, but it will not display its label.

x = y + z  (6)

If you want to typeset several columns of equations, it is quite easy:

x = y + z  a = b + c  x = v  (7)
x = y + z  a = b + c  x = u + 1  (7’)
(multicol)  x = y
a = b  (multicol)
x = y + z  a^{2} = (b − c)^{3} + y

and also (this is some \intertext)

x = y + z  a = (b + c)^{2} − 5  ℓ = m  (8)

You can even extend the block (5) much later like

x = y + z x = y + z f(x) = b  (5d)
x = y + z x = y + z g(x) = b  (5e)
sin^{2} x + cos^{2} x = 1  (5f)

You can also change the type of the subnumbering and use the \text command without loading amstext. Like for example

I_{1} = ∫_{a}^{b} g(x) dx ,  (First equation)  (9A)
I_{2} = ∫_{a}^{b} g(x^{2} − 1) dx . (Second equation)  (9B)